# Building An Open Architecture for AI-RAN Convergence in 6G

### v1

@@ Abstract
The proliferation of data-intensive Artificial Intelligence (AI) applications at the network edge demands a fundamental shift in RAN design, from merely consuming AI for network optimization, to actively enabling distributed AI workloads. This paradigm shift presents a significant opportunity for network operators to monetize AI at the edge while leveraging existing infrastructure investments. To realize this vision, this article presents a novel converged O-RAN and AI-RAN architecture that unifies orchestration and management of both telecommunications and AI workloads on shared infrastructure.
The proposed architecture extends the Open RAN principles of modularity, disaggregation, and cloud-nativeness to support heterogeneous AI deployments. We introduce two key architectural innovations: (i) the AI-RAN Orchestrator, which extends the O-RAN Service Management and Orchestration (SMO) to enable integrated resource and allocation across RAN and AI workloads; and (ii) AI-RAN sites that provide distributed edge AI platforms with real-time processing capabilities. The proposed system supports flexible deployment options, allowing AI workloads to be orchestrated with specific timing requirements (real-time or batch processing) and geographic targeting.
The proposed architecture addresses the orchestration requirements for managing heterogeneous workloads at different time scales while maintaining open, standardized interfaces and multi-vendor interoperability.
@@ I Introduction
5G-Advanced and 6G networks are reshaping the role of the RAN, from a traditional communication substrate to a programmable, intelligent, “sentient”, and compute-capable platform. As data-intensive and latency-sensitive applications such as Extended Reality (XR), Integrated Sensing and Communications (ISAC), and generative Artificial Intelligence (AI) emerge, there is an opportunity to rethink RAN architecture not only as a consumer of AI for internal optimization, but also as a distributed platform for hosting AI workloads [1]. This transformation demands architectural innovation that supports the convergence of communication and AI computing within the network itself—simultaneously delivering connectivity and supporting distributed AI inference and training demands. This dual capability is a foundational principle of the AI-RAN Alliance, a new telecom industry initiative dedicated to transforming networks through pervasive AI integration [2].
@@
This paradigm shift in telecom infrastructure demands new approaches to system architecture and management. In this paper, we design network resource orchestration and deployment workflows to support AI solutions (i) for the RAN, i.e., tasked with improving its efficiency and operations, or (ii) on the RAN, i.e., leveraging an extremely distributed compute infrastructure for inference, sensing, or training at the edge. The network infrastructure and orchestration abstraction layer supports flexibility, scalability, and interoperability in enabling innovative AI-native wireless systems for next-generation applications. At the same time, to enable innovation, this network abstraction and the underlying infrastructure need to embrace openness, standardized interfaces, disaggregated components and support for multi-vendor deployments.
Figure 1: Architecture enabling AI and RAN coexistence, integrated within the proposed extension of the O-RAN framework.
@@
Despite preliminary efforts in edge AI deployment, the telecom ecosystem currently lacks a cohesive architectural framework that can effectively manage the complex interplay between AI workloads and RAN operations at scale. The O-RAN ALLIANCE is defining open, interoperable interfaces and introducing key components such as the Service Management and Orchestration (SMO) and the RAN Intelligent Controllers (RICs), which enable intelligent control across multiple latency domains [3]. While O-RAN leverages intelligent control loops to optimize RAN performance, its architecture does not support generic AI workloads through a shared RAN infrastructure. As discussed above, this is currently being considered by the AI-RAN Alliance, which is defining coexistence mechanisms across different workloads. Recent literature has also addressed AI and RAN orchestration [4, 5, 1, 6, 7]. Specifically, [4, 5] discuss AI operations in the cloud, edge, and also RAN, orchestrating communications, computing, data, and AI models. [6] proposes a converged AI-and-O-RAN architecture with an end-to-end orchestrator that interacts with the Near-RT RIC its programmable logic units (xApps), thus limiting the scope of services that can be managed and deployed. [7] proposes operational workflows for AI-RAN. [1] presents a converged compute-communication platform that integrates both RAN and AI workloads on the same physical infrastructure. However, a comprehensive, end-to-end, and generalized architecture enabling AI and RAN coexistence following O-RAN management and orchestration principles is still missing.
@@
Contributions. This paper takes a decisive step in this direction by proposing a converged O-RAN and AI-RAN architecture, where the O-RAN framework is extended to support the orchestration and management of AI workloads, as shown in Fig. 1. By building upon the foundational principles of modularity, disaggregation, and cloud-nativeness of the existing O-RAN architecture, this work demonstrates how AI integration can be achieved with minimal disruption and without increasing system complexity. The proposed design extends that the RAN compute infrastructure to the AI datacenter domain, introducing new monetization opportunities for public and private operators. Specifically, we make the following key contributions:
@@
$\bullet$ Gap Analysis. We review AI-RAN activities and discuss the technical and operational challenges that must be addressed to achieve seamless AI-RAN and O-RAN integration.
@@
$\bullet$ Unified Architecture Design. We present a new comprehensive architecture that enables AI and RAN workloads to share infrastructure resources while maintaining performance isolation and service quality, shown in Fig. 1. We also identify and address the orchestration requirements for managing heterogeneous AI and RAN workloads at different time scales. Specifically, we identify and propose two key architectural components, (i) the AI-RAN Orchestrator, at the top of Fig. 1, extends the O-RAN SMO framework to provide unified resource allocation, policy management, and workload coordination across AI and RAN services; and (ii) the AI-RAN Sites, which augment the O-RAN O-Cloud concept to further support AI within a common RAN infrastructure, enabling a distributed edge AI platform with real-time processing capabilities.
@@
$\bullet$ Flexible Orchestration. We define comprehensive procedures for deploying AI workloads on shared RAN infrastructure, introducing two distinct workflows: (i) Batch mode for non-real-time computing requests and analytical workloads that are flexible with respect to deployment site and timeline; and (ii) Real-time mode for latency-critical AI workloads that require immediate processing and response on specific AI-RAN sites.
@@ II Toward AI-RAN
The transition towards AI-native mobile networks is being driven by three complementary research tracks within the AI-RAN Alliance: AI-for-RAN, AI-on-RAN, and AI-and-RAN. This section reviews the key principles behind each AI-RAN working group. We then analyzes how these efforts align with O-RAN ALLIANCE specification activities to establish the technical foundation for our converged architecture presented in Sec. IV.
@@ II-A AI-RAN Principles
AI-for-RAN. The first pillar, AI-for-RAN, applies AI to enhance network automation, efficiency, and adaptability [5]. AI-for-RAN enhances RAN management through data-driven decision-making, improving performance, reliability, and cost efficiency via automation and self-healing [8, 2]. While this enables intelligent, self-optimizing networks, its scope remains limited to optimizing existing RAN functions without extending AI’s role beyond network operations [2, 3].
@@
AI-on-RAN. Conversely, AI-on-RAN enables AI and generative AI applications to execute on the RAN infrastructure. These applications may or may not be related to cellular systems, but can leverage spare compute capacity and the geographically distributed nature of the RAN infrastructure. Traditional RAN deployments are over-provisioned for peak loads, therefore, by leveraging AI during low RAN activity periods, operators can maximize infrastructure efficiency without compromising real-time cellular communication [4]. This allows operators to provide AI-as-a-Service for diverse AI-driven applications while ensuring high-performance edge computing. By embedding AI workloads within the RAN, AI-on-RAN brings AI close to end-users and thus reduces the latency and bandwidth overhead associated with sending data to centralized data centers, making it ideal for latency-sensitive applications that require real-time inference and decision making. Examples of AI-on-RAN workloads can include distributed AI training (e.g., federated learning), video processing and computer vision applications, or edge deployments of generative AI solutions (e.g., a low-latency edge instance of an Large Language Model (LLM)), among others.
@@
AI-and-RAN. To enable infrastructure sharing across AI applications and RAN workloads, AI-and-RAN focuses on the coexistence of heterogeneous workloads on the same compute platform. Indeed, AI and RAN workloads differ in terms of reliability, availability, and performance requirements. The RAN digital signal processing represents a non-elastic workload, where compute tasks need to be executed within specific timing constraints to ensure that the overall protocol stack can function as intended (e.g., for channel coding/decoding, scheduling, among others). At the same time, processing for a specific carrier configuration (bandwidth, number of MIMO streams) requires a fixed and predictable amount of resources. AI and generative AI inference or training solutions are more elastic, provided that enough resources are allocated to load the data structures associated with the AI models.
Figure 2: Challenges toward the integration of AI and RAN within the O-RAN architecture.
@@
AI-RAN Orchestration Requirements. Considering these requirements, the orchestration of AI-for/on-RAN requires (i) a holistic accounting of the underlying compute and networking resources, so as to dynamically track infrastructure availability; (ii) northbound Application Programming Interfaces (APIs) to AI and RAN developers, to expose available capacity with a high granularity and to enable submitting workloads to the system, thus enabling infrastructure monetization; (iii) an interface that allows operators to express requirements and intents for the AI-RAN orchestrator (e.g., priority level for RAN workloads, maximum latency for AI inference); (iv) a scheduling strategy that combines requests, intents, and available resources into actionable policies or allocations; and (v) a southbound interface that connects to the infrastructure for AI and RAN configuration and deployment, and also allows for preemption in case time-sensitive RAN workloads need to be prioritized. Figure 1 captures these requirements in the end-to-end architecture.
II-B The Role of Open RAN
The AI-RAN Alliance mission does not include specifications or standards development. As such, it operates considering technical developments, specifications and standardization efforts within 3GPP and the O-RAN ALLIANCE. The latter has driven the design and development of networked systems that abide by the principles of openness, programmability, interoperability, and intelligent closed-loop control [3]. While focusing primarily on AI-for-RAN, leveraging AI to optimize RAN performance, the O-RAN ALLIANCE architecture has introduced several elements that provide valuable opportunities to extend AI integration beyond optimization. These include a software-based, disaggregated RAN, with atomic network functions that are deployed on the O-Cloud, a virtualization platform for cellular systems; the RICs, which connect to the RAN through open interfaces and enable the onboarding of custom intelligence in xApps (for the Near-RT RIC) and rApps (for the Non-RT RIC); the SMO, which provides management and orchestration across the whole network; and pipelines and workflows for AI-for-RAN.
@@ II-B The Role of Open RAN
The SMO and Non-RT RIC concern network management in the time scale of seconds or more, for RAN services and RAN performance/efficiency, respectively. The latest O-RAN architecture decouples them in a service-based framework, where the SMO combines multiple independent services communicating through the standardized SMO Services (SMOS) interface. The O2 interface connects the SMO with the O-Cloud [9] for infrastructure management and workload deployment.
We argue that the cloud-native architecture of Open RAN can streamline the development and adoption of AI-RAN innovation, through open interfaces and service composition that allow to onboard software, workloads, and workflows from multiple vendors, providers, and users. The open ecosystem fosters multi-vendor collaboration, ensuring that AI algorithms can be developed and deployed across diverse hardware and software platforms without vendor lock-in and with well-defined procedures. In addition, building on an existing architecture makes it possible to avoid unnecessary complexity, extending O-RAN components to also manage AI workloads rather than introducing additional orchestration solutions.
Building on this foundation, the following section identifies the critical challenges facing converged open AI-RAN architectures and how to address them.
@@ III Challenges and Open Issues Toward AI-RAN and O-RAN Convergence
Figure 2 summarizes the main challenges toward effective AI-and-RAN orchestration and integration with O-RAN systems. These challenges arise from deploying AI into O-RAN ’s open and programmable architecture, which requires robust coordination between AI and O-RAN-compliant network functions.
@@
Architecture and Workflows Challenges. AI-RAN coexistence requires an evolution of the 3GPP and O-RAN architectures. In the control plane, the end-to-end management system needs to include an entity responsible for the coexistence between AI and RAN, which maps the functionalities of the orchestrator introduced in Fig. 1 and in Sec. II. This includes understanding what are the proper control and data exposure abstractions for such a task, what is the component of the network that can host such functionalities, and if there needs to be a hierarchical approach to orchestration.
@@
From a workflow point of view, there is a need for data, telemetry, control, and deployment pipelines, extending the O-RAN interfaces. This includes the definition of the southbound and northbound interfaces of the orchestrator, as shown in Fig. 1 at a high level. Southbound interfaces should support the continuous collection, pre-processing, and delivery of high-quality telemetry from distributed AI-RAN components, accommodating the latency requirements of different monitoring, inference, or control applications. The design of these interfaces comes with the unique opportunity to enable open and systematic processes for the creation of large-scale datasets. Finally, what Key Performance Measurements (KPMs) are better suited to support orchestrator tasks is an open question.
@@
AI-RAN-specific procedures must be defined within the architectural framework to manage the lifecycle of AI workloads. This implies regulating access to RAN resources, orchestrating deployment across nodes, and ensuring compliance with the operator’s policies. To do so, current O-RAN interfaces (e.g., O1, O2, E2, A1) require targeted modifications to support the dual operation of RAN control loops and AI workflows. Without these foundational elements, the AI-RAN vision cannot scale beyond isolated use cases into a fully integrated, intelligent network infrastructure.
@@
Orchestration Challenges. From an infrastructure point of view, as shown in Fig. 2, an AI-RAN system includes a heterogeneous pool that combines compute with different AI accelerators (e.g., Graphics Processing Units (GPUs)), storage, networking elements, radios, and spectrum, which the orchestrator needs to jointly take into account when allocating resources. RAN nodes exhibit diverse processing capabilities, also based on deployment location. Differences in computational power, memory, and latency requirements can lead to inefficiencies in workload distribution [1, 10]. The heterogeneity also extends to the workloads: as discussed in Sec. II, AI and RAN often have different workload patterns and requirements. AI itself can range from lightweight inference for real-time network optimization or value-added services to high-intensity model training. Moreover, the unpredictability of AI workloads, and in particular that of large-scale training tasks, can introduce latency spikes that may negatively impact RAN performance [11, 12]. This complicates both monitoring, which needs to account for real-time variations in the workload intensity, and orchestration, which needs to combine multi-scale and proactive coexistence approaches.
@@
Resource Allocation Challenges. The resource allocation problem for the orchestrator needs to keep into account multiple factors including operators’ priorities, the multi-dimensionality of the problem (encompassing a distributed infrastructure across different geographical locations with heterogeneous capabilities), and the potential presence of a hierarchical control structure with varying granularity, from a centralized orchestrator to distributed components. The orchestrator input also comes from different stakeholders (e.g., AI developers, RAN vendors, operators), whose orthogonal requirements may need to be harmonized and reconciled.
@@
Use Case Challenges. The AI-RAN orchestrator enables monetization strategies for next-generation wireless. These depend on identifying trade-offs, prioritization, and provisioning strategies to satisfy both AI and RAN tenants. To successfully cater to stakeholders representing AI-on-RAN use cases, the orchestrator northbound AI APIs need to align with best practices in the AI and cloud industries. The coexistence and monetization challenge also extends to energy efficiency. With RAN energy consumption being the highest operational expenditure for operators to date, it is important to adopt an energy-efficient and opportunistic approach to the acceleration of AI. AI models are usually associated with high computational demands, in contrast to recent trends for low-power RAN components. Hardware acceleration offers potential mitigation but requires careful scheduling to ensure energy-efficient performance [11].
@@
Integrating third-party AI workloads also presents challenges to security, privacy, and access management. The orchestrator APIs and the end-to-end control and infrastructure management pipelines need to ensure accountability and tracking of security roles and privileges. One of the primary concerns is the deployment of AI models or workloads which include vulnerabilities in the software stack that may affect the stability of the RAN. Given the critical nature of RAN operations, such attacks could lead to network congestion, degraded service quality, or even large-scale outages [13].

### v2

# Beyond Connectivity: An Open Architecture for AI-RAN Convergence in 6G

@@ Abstract
Data-intensive Artificial Intelligence (AI) applications at the network edge demand a fundamental shift in RAN design, from merely consuming AI for network optimization, to actively enabling distributed AI workloads. This presents a significant opportunity for network operators to monetize AI while leveraging existing infrastructure. To realize this vision, this article presents a novel converged O-RAN and AI-RAN architecture for unified orchestration and management of telecommunications and AI workloads on shared infrastructure.
The proposed architecture extends the Open RAN principles of modularity, disaggregation, and cloud-nativeness to support heterogeneous AI deployments. We introduce two key architectural innovations: (i) the AI-RAN Orchestrator, which extends the O-RAN Service Management and Orchestration (SMO) to enable integrated resource and allocation across RAN and AI workloads; and (ii) AI-RAN sites that provide distributed edge AI platforms with real-time processing capabilities. The proposed architecture enables flexible orchestration, meeting requirements for managing heterogeneous workloads at different time scales while maintaining open, standardized interfaces and multi-vendor interoperability.
@@ I Introduction
5G-Advanced and 6G networks are reshaping the role of the RAN, from a traditional communication substrate to a programmable, intelligent, “sentient”, and compute-capable platform. With emerging data-intensive and latency-sensitive use cases such as Extended Reality (XR), Integrated Sensing and Communications (ISAC), and generative AI, there is an opportunity to rethink RAN architecture not only as a consumer of AI for internal optimization, but also as a distributed platform for hosting AI workloads [1]. This transformation demands architectural innovation that supports the convergence of communication and AI computing within the network itself—simultaneously delivering connectivity and supporting distributed AI inference and training demands. This dual capability is a foundational principle of the AI-RAN Alliance, a new telecom industry initiative dedicated to transforming networks through pervasive AI integration [2].
@@
In this paper, we design network resource orchestration and deployment workflows to enable infrastructure support for AI solutions (i) for the RAN, i.e., tasked with improving efficiency and operations; or (ii) on the RAN, i.e., leveraging the distributed compute infrastructure for inference, sensing, or training at the edge. The network infrastructure and orchestration abstraction layer supports flexibility, scalability, and interoperability to enable innovative AI-native next-generation wireless systems. At the same time, to enable innovation, this network abstraction and the underlying infrastructure need to embrace openness, standardized interfaces, disaggregated components and support for multi-vendor deployments.
Figure 1: Architecture enabling AI and RAN coexistence, integrated within the proposed extension of the O-RAN framework. The figure also includes an example of use case for AI-and-RAN coexistence. An application developer submits an AI-on-RAN workload to enable XR processing at the edge. In the meantime, a RAN vendor provides an updated AI-for-RAN slicing solution. The operator defines the policy that the infrastructure should follow to enable both services. The AI-RAN orchestrator translates the policy into resource allocation, and coordinates with the automation workflows to deploy the services in the AI-RAN sites.
@@
In this area, the prior work focused on edge AI, open architectures, and early results on AI and RAN orchestration. Preliminary efforts in edge AI deployment [3] did not lead to a cohesive architectural framework that can effectively manage the complex interplay between AI workloads and RAN operations at scale. The O-RAN ALLIANCE introduced intelligent control loops to optimize RAN performance [4]. However, its architecture, with open, interoperable interfaces, a SMO and the RAN Intelligent Controllers (RICs), does not support generic AI workloads on a shared RAN infrastructure. As discussed above, this is instead a focus area for the AI-RAN Alliance. Recent literature has also addressed AI and RAN orchestration [5, 6, 1, 7, 8]. Specifically, [5] and [6] discuss AI operations in the cloud, edge, and also RAN, but do not propose an architecture for orchestration. [7] proposes a converged AI-and-O-RAN architecture but the end-to-end orchestrator only interacts with the Near-RT RIC programmable logic units (xApps), thus limiting the scope of services that can be managed and deployed. [8] discusses AI-RAN AI operational workflows for AI-RAN, but does not consider architectural elements. [1] presents an infrastructure design that jointly supports RAN and AI workloads, without considering AI-RAN orchestration. Therefore, a comprehensive, end-to-end, and generalized architecture enabling AI and RAN coexistence following O-RAN management and orchestration principles is still missing.
@@
Contributions. This paper takes a decisive step in this direction by proposing a converged O-RAN and AI-RAN architecture, where the O-RAN framework is extended to support the orchestration and management of AI workloads, as shown in Fig. 1. By building upon the foundational principles of modularity, disaggregation, and cloud-nativeness of the existing O-RAN architecture, this work demonstrates how AI integration can be achieved with minimal disruption and without increasing system complexity. The proposed design extends the RAN compute infrastructure to the AI datacenter domain, introducing new monetization opportunities for public and private operators. Specifically, we make the following key contributions:
@@
$\bullet$ Gap Analysis. We review AI-RAN activities and discuss the technical and operational challenges to achieve seamless AI-RAN and O-RAN integration.
@@
$\bullet$ Unified Architecture Design. We present a new comprehensive architecture that enables AI and RAN workloads to share infrastructure while maintaining performance isolation and service quality, shown in Fig. 1. We also identify and address the orchestration requirements for managing heterogeneous AI and RAN workloads at different time scales. Specifically, we identify and propose two key architectural components. (i) The AI-RAN Orchestrator (Fig. 1, top) extends the O-RAN SMO framework for unified resource allocation, policy management, and workload coordination across AI and RAN services. (ii) The AI-RAN Sites augment the O-RAN O-Cloud concept to further support AI within a common RAN infrastructure, enabling a distributed edge AI platform with real-time processing capabilities.
@@
$\bullet$ Flexible Orchestration. We define comprehensive procedures for deploying AI workloads on shared RAN infrastructure, introducing two distinct workflows for non-real-time batch and latency-critical deployments.
@@ II Toward AI-RAN
This section reviews the key elements of each AI-RAN Alliance working group (AI-for-RAN, AI-on-RAN, AI-and-RAN) and the O-RAN ALLIANCE architecture.
@@ II-A AI-RAN Principles
AI-for-RAN. The first pillar, AI-for-RAN, applies AI to enhance network automation, efficiency, and adaptability [6]. AI-for-RAN enhances RAN management through data-driven decision-making and Digital Signal Processing (DSP), improving performance, reliability, and cost efficiency [2]. This enables intelligence and self-optimization in existing RAN functions [2, 4].
@@
AI-on-RAN. Conversely, AI-on-RAN enables AI and generative AI applications to execute on the RAN infrastructure. These applications may or may not be related to cellular systems, but leverage spare compute capacity and the geographically distributed nature of the RAN. RAN deployments are over-provisioned for peak loads, therefore, by leveraging AI during low RAN activity periods, operators can maximize infrastructure efficiency without compromising connectivity [5]. By embedding AI workloads within the RAN, AI-on-RAN brings AI close to end-users and thus reduces the latency and bandwidth overhead for transmission to centralized data centers. This makes AI-on-RAN ideal for latency-sensitive applications that require real-time inference and decision making. It also helps distributing workloads from large data centers to the edge. Examples of AI-on-RAN workloads can include distributed AI training (e.g., federated learning), video processing and computer vision applications, or edge generative AI solutions (e.g., a low-latency edge instance of a Large Language Model (LLM), chatbots, AI agents), among others.
@@
AI-and-RAN. To enable infrastructure sharing across AI applications and RAN workloads, AI-and-RAN focuses on the coexistence of heterogeneous workloads. Indeed, AI and RAN workloads differ in terms of reliability, availability, and performance requirements. The RAN DSP represents a non-elastic workload, where tasks need to be executed within specific timing constraints to ensure intended functionality of the protocol stack (e.g., for channel coding/decoding, scheduling). At the same time, processing for a specific configuration (bandwidth, number of MIMO streams) requires a fixed and predictable amount of resources. AI and generative AI inference or training solutions are more elastic, provided that enough resources are allocated to load the data structures associated with the AI models.
@@
AI-RAN Orchestration Requirements. Considering these requirements, the orchestration of AI-for/on-RAN requires (i) a holistic accounting of the underlying compute and networking resources, so as to dynamically track infrastructure availability; (ii) northbound Application Programming Interfaces (APIs) to AI and RAN developers, to expose available capacity and to submit workloads to the system, thus enabling infrastructure monetization; (iii) an interface for operator requirements/intents (e.g., priority level for RAN workloads, maximum latency for AI inference); (iv) a scheduling strategy that combines requests, intents, and available resources into actionable policies or allocations; and (v) a southbound interface for AI and RAN configuration and deployment. This also allows for preemption for time-sensitive RAN workloads. Figure 1 captures these requirements in the end-to-end architecture.
II-B A Primer on Open RAN
The AI-RAN Alliance does not develop specifications or standards, but builds on and influences those within 3GPP and the O-RAN ALLIANCE. The latter has driven the design and development of networked systems that abide by the principles of openness, programmability, interoperability, and intelligent closed-loop control. A complete overview can be found in [4].
The O-RAN architecture is based on a softwarized, disaggregated RAN, with atomic network functions that are deployed on the O-Cloud, a virtualization platform for cellular systems. Through open interfaces, the RICs connect the RAN to custom intelligence in xApps (for the Near-RT RIC) and rApps (for the Non-RT RIC). The SMO provides management and orchestration across the whole network. The O-RAN architecture already features pipelines and workflows for AI-for-RAN.
@@ II-B A Primer on Open RAN
The SMO and Non-RT RIC concern network management in the time scale of seconds or more, for RAN services and RAN performance/efficiency, respectively. The latest O-RAN architecture decouples them in a service-based framework, where the SMO combines multiple independent services communicating through the standardized SMO Services (SMOS) interface. The O2 interface connects the SMO with the O-Cloud for infrastructure management and workload deployment [9].
@@ III Challenges and Open Issues Toward AI-RAN and O-RAN Convergence
The cloud-native architecture of Open RAN can streamline AI-RAN innovation. Open interfaces and service composition allow onboarding software, workloads, and workflows from multiple vendors, providers, and users. The open ecosystem fosters multi-vendor collaboration, ensuring that AI algorithms can be developed and deployed across diverse hardware and software platforms without vendor lock-in and with well-defined procedures. In addition, building on an existing architecture makes it possible to avoid unnecessary complexity, extending O-RAN components to also manage AI workloads rather than introducing additional orchestration solutions.
Effective AI-and-RAN orchestration and integration with O-RAN system, however, requires addressing several challenges, which we discuss next.
@@
Architecture and Orchestration Workflows Challenges. AI-RAN coexistence requires an evolution of the 3GPP and O-RAN architectures. The end-to-end management system requires an entity responsible for AI-and-RAN coexistence, i.e., the orchestrator introduced in Fig. 1 and in Sec. II. Challenges include understanding what are effective control and data exposure abstractions, which is network host for such functionalities, and if orchestration requires a hierarchical approach.
@@
From a workflow point of view, there is a need for telemetry, control, and deployment pipelines, extending the O-RAN interfaces to define the southbound and northbound interfaces of the orchestrator (Fig. 1). Southbound interfaces should support continuous telemetry collection, preprocessing, and delivery from distributed AI-RAN components, accommodating latency requirements for monitoring, inference, or control applications. This unlocks the unique opportunity to systematically create large-scale datasets. Finally, what Key Performance Measurements (KPMs) are better suited to support orchestrator tasks is an open question.
@@
Finally, the definition of AI-RAN-specific procedures for AI lifecycle management implies regulating access to RAN resources, orchestrating deployment across nodes, and ensuring compliance with the operator’s policies. To do so, current O-RAN interfaces (e.g., O1, O2, E2, A1) require targeted modifications to support the dual operation of RAN control loops and AI workflows.
@@
Orchestration Challenges. An AI-RAN system (as shown in Fig. 1) combines compute with different AI accelerators (e.g., Graphics Processing Units (GPUs)), storage, networking elements, radios, and spectrum. The orchestrator needs to account for such heterogeneity when allocating resources. Additional compute can be available in the core network, which however does not have the same distributed nature as the RAN. RAN nodes exhibit diverse processing capabilities, also based on deployment location, which can lead to inefficiencies in workload distribution [1, 10]. The heterogeneity also extends to the workloads. As discussed in Sec. II, AI and RAN often have different workload patterns and requirements. AI itself can range from lightweight inference for real-time network optimization or value-added services to high-intensity model training. Moreover, the unpredictability of AI workloads, and in particular that of large-scale training tasks, can introduce latency spikes that may negatively impact RAN performance [11, 12]. This complicates both monitoring, which needs to account for real-time variations in the workload intensity, and orchestration, which needs to combine multi-scale and proactive coexistence approaches. Finally, orchestration workflows can extend beyond the RAN, toward core network functions that contribute to the AI operations.
@@
Resource Allocation Challenges. The resource allocation problem must consider operators’ priorities, the multi-dimensionality of the problem (encompassing a distributed infrastructure across different geographical locations with heterogeneous capabilities), and potentially a hierarchical control structure with varying granularity, from a centralized orchestrator to distributed components. The orchestrator input also comes from different stakeholders (e.g., AI developers, RAN vendors, operators), whose orthogonal requirements may need to be harmonized and reconciled. Moreover, operators may impose different Service Level Agreement (SLA) to AI and RAN. These may change across sites and time and need to be accounted for when allocating resources. Therefore, while solutions can be inspired by prior work on scheduling and slicing, they also need to take into account the complexity of the input, output, and dimensionality of the infrastructure partitioning problem.
Use Case, Monetization, and Energy Efficiency Challenges. The AI-RAN orchestrator enables monetization strategies for next-generation wireless by identifying trade-offs, prioritizing, and provisioning strategies to satisfy both AI and RAN tenants. To successfully cater to AI-on-RAN stakeholders, the orchestrator needs to align with AI and cloud best practices.
@@
The coexistence and monetization challenge also extends to energy efficiency. With RAN energy consumption being the highest operational expenditure for operators to date, it is important to adopt an energy-efficient and opportunistic approach to the acceleration of AI, and to account for the impact of energy in the total costs of the AI-RAN infrastructure. AI models are usually associated with high computational demands, in contrast to recent trends for low-power RAN components. Hardware acceleration offers potential mitigation but requires careful scheduling to ensure energy efficiency [11] while accounting for non-linearity and multi-dimensionality of the AI accelerators power consumption as a function of processor load, memory utilization, and data transfer.
@@
Integrating third-party AI workloads also presents challenges to security, privacy, and access management. The orchestrator APIs and the end-to-end control and infrastructure management pipelines need to ensure accountability and tracking of security roles and privileges. Deploying AI models or workloads with vulnerabilities in the software stack could affect the stability of the RAN. Such attacks could lead to network congestion, degraded service quality, or large-scale outages [13].